# Codes Correcting Two Bursts of Exactly $b$ Deletions

### v4

@@ II-A Related Results
Recently, there have been two notable developments in non-binary two-deletion correcting codes with low redundancy. In [21, Theorem 1], the authors presented a $q$-ary two-deletion correcting code with redundancy at most $5\log n+(16\log q+10)\log\log n+o(\log\log n)$, for all even $q>2$. In [22], a $q$-ary two-deletion correcting code was constructed, with redundancy at most $5\log n+10\log\log n+O_{q}(1)$ (where $O_{q}(1)$ denotes a constant depending only on $q$), for all $q>2$. In Section V, we present a new construction of $q$-ary two-deletion correcting codes with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$, for all $q>2$.
Regarding codes that can correct two $b$-burst-deletions (where $b>1$), there are, to our knowledge, no explicit results available. A related result can be found in the work of Sima et al [33, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ with the deletions in each burst not necessarily occurring consecutively. For $t=2$ and $t_{L}=b$, their result provides a binary $2$-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$. Their result was derived using the syndrome compression technique, which was later extended to syndrome compression with pre-coding in [20]. We will apply this extended technique to provide a $q$-ary code with redundancy at most $7\log n+O\left(\log n/\log\log n\right)$, for all $q\geq 2$.
@@ 1. $(\textup{{P}}1)$
Then there exists a function $\bar{f}:\mathcal{E}\rightarrow\{0,1\}^{2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)}$, computable in polynomial time, such that $\bar{f}\left(\mathbf{x}\right)\neq\bar{f}\left(\mathbf{y}\right)$ for any $\mathbf{x}\in\mathcal{E}$ and $\mathbf{y}\in\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)$.
Let $\mathcal{E}$ be a $1$-$b$-burst correcting code and $\bar{f}$ be given in Lemma II.1. Then Lemma II.1 asserts that if $\mathbf{x},\mathbf{y}\in\mathcal{E}$ are distinct codewords and $\bar{f}(\mathbf{x})=\bar{f}(\mathbf{y})$, we have $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})=\emptyset$. Therefore, for any $\mathbf{a}\in\{0,1\}^{2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)}$, the code =$\mathcal{E}^{\prime}=\left\{\mathbf{x}\in\mathcal{E}:\bar{f}(\mathbf{x})=a\right\}$ is a $2$-$b$-burst-deletion correcting code. Furthermore, by the pigeonhole principle, there exists an $\mathbf{a}$ such that the redundancy of $\mathcal{E}^{\prime}$ is at most $\rho\left(\mathcal{E}^{\prime}\right)=\rho(\mathcal{E})+2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)$ [^2] We select $\mathcal{E}$ to be a $1$-$b$-burst correcting code to obtain a better redundancy. If we take $\mathcal{E}=\Sigma_{q}^{n}$ we get a redundancy of $8\log n+o(\log n)$..
@@
[32, Theorem 9, $t=b,s=0$] For all $q\geq 2$ and $n>b$, there is a function $\phi:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{\log n+O_{q,b}(1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\phi(\mathbf{x})$ and $\mathbf{y}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$. Here, $O_{q,b}(1)$ is a constant dependent only on $q$ and $b$.
This lemma gives a $1$-$b$-burst-deletion correcting code $\mathcal{E}$ with redundancy $\log n+O_{q,b}(1)$. Since $\mathcal{E}$ can correct single $b$-burst-deletion, by a simple counting, we can see that $\left|\mathcal{N}_{\mathcal{E}}(\mathbf{x})\right|\leq q^{b}n^{3}$. In fact, each codeword in $\mathcal{N}_{\mathcal{E}}(\mathbf{x})$ can be obtained in the following three steps:
@@
    Delete two substrings of length $b$ from $\mathbf{x}$, resulting in $\mathbf{z}^{(1)}$. There are at most $n^{2}$ possibilities for $\mathbf{z}^{(1)}$.
@@
    Insert a sequence of length $b$ into $\mathbf{z}^{(2)}$ to get a sequence $\mathbf{y}\in\Sigma_{q}^{n}$. Since $\mathcal{E}$ is a $1$-$b$-burst-deletion correcting code, for each $\mathbf{z}^{(2)}$, there is at most one $\mathbf{y}$ which is in $\mathcal{N}_{\mathcal{E}}(\mathbf{x})$.
@@
[17, Theorem 2] For any integer $n\geq 3$, there exists a function $\xi:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{7\log n+O\left(1\right)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{2}^{n}$, given $\xi\left(\mathbf{x}\right)$ and any $\mathbf{y}\in\mathcal{B}_{1}^{2}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$.
@@
where $x_{k,i}\in\{0,1\}$ be such that $x_{i}=\sum_{k=1}^{\left\lceil\log q\right\rceil}x_{k,i}2^{k-1}$ for all $1\leq i\leq n$.
@@
Let $A(\mathbf{x})=A(\mathbf{x},n,b)$ and $A(\mathbf{y})=A(\mathbf{y},n-b,b)$. Since $\mathbf{y}\in\mathcal{B}_{2}^{b}(\mathbf{x})$, we have $A(\mathbf{y})_{i}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$. Let $\xi_{1}(\cdot)$ be the function defined in Lemma II.4. Define $\psi(\mathbf{x})\triangleq\left(\xi_{1}\left(A(\mathbf{x})_{1}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b}\right)\right)$. Then by Lemma II.4, given $\mathbf{y}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$ and $\psi(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$. Since each $\xi_{1}(A(\mathbf{x})_{i})$ is a binary vector of length $7\left\lceil\log q\right\rceil\log(n/b)+O_{q}(1)$, we can see that $\psi(\mathbf{x})$ is a binary vector of length $7b\left\lceil\log q\right\rceil\log(n/b)+O_{q,b}(1)$. ∎
@@
Before proceeding to subsequent sections, we introduce two useful lemmas. The first one is a well-known result from number theory. The second one will be used in Section IV-B.
Lemma II.6 (Bertrand’s Postulate/Bertrand–Chebyshev theorem)
Let $N\geq 2$ be an integer. Suppose $\mathbf{x}\in[0,N-1]^{n}$, where $n>4$. Let ???? denote an unknown symbol (not in $[0,N-1]$). If $\mathbf{y}\in\left([0,N-1]\cup\{?\}\right)^{n}$ be such that $y_{i}=y_{i+1}=y_{j}=y_{j+1}=?$ and $y_{k}=x_{k}$ for all $k\notin\{i,i+1,j,j+1\}$, we say that $\mathbf{y}$ is obtained from $\mathbf{x}$ by two bursts of erasures (of length two). For our purpose in Section IV-B, assume $j\geq i+2$. For a sequence $\mathbf{x}$ over the alphabet $[0,N-1]$, denote $A(\mathbf{x})=A(\mathbf{x},\left|\mathbf{x}\right|,2)$ and $\textup{Syn}\left(\mathbf{x}\right)\triangleq\sum_{i=1}^{n}ix_{i}$.
Lemma II.7
For any $0\leq a_{1},a_{2}<2N$ and $0\leq b<nN^{2}$, define $\mathcal{C}$ to be set of all sequences $\mathbf{x}\in[0,N-1]^{n}$ that satisfies:
@@
$\varphi(\mathbf{x})\triangleq\left(\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{1,j}\pmod{2N},\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2,j}\pmod{2N},W(\mathbf{x})\pmod{nN^{2}}\right).$
@@
$\Delta=\left(b-\sum_{j=1,j\neq i_{1},i_{2}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{y})_{1,j}-(2N-1)\sum_{j=1,j\neq i_{3},i_{4}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{y})_{2,j}\right)\pmod{nN^{2}}.$ (2)
Since each term in the right-hand side of (2) is known, we can obtain the value of $\Delta$. Furthermore, by (2), Condition (C2), and the relationship between $A(\mathbf{x})$ and $A(\mathbf{y})$, it holds that
@@
Since $i_{1}<i_{2}\leq\left\lceil n/2\right\rceil$ and $i_{4}-i_{3}\leq\left\lceil n/2\right\rceil$, we have $21\leq i_{2}-i_{1},i_{4}-i_{3}\leq n/$ 2. Combining this with the fact $10\leq\alpha_{1},\alpha_{2}\leq N-$ 1, we conclude that $0\leq(i_{2}-i_{1})\alpha_{1}+(2N-1)(i_{4}-i_{3})\alpha_{2}<nN^{2}$. Now Equation 4 implies
@@
Note that $10\leq\alpha_{1}\leq N-$ 1. When $i_{4}-i_{3}=i_{2}-i_{1}$ or $i_{2}-i_{1}+1$, it it easy to see that $0\leq(i_{2}-i_{1})\alpha_{1}<(2N-1)(i_{4}-i_{3})$. When $i_{4}-i_{3}=i_{2}-i_{1}-1$, since $i_{2}-i_{1}\geq 2$, we have $(2N-1)(i_{4}-i_{3})-(i_{2}-i_{1})\alpha_{1}=(i_{2}-i_{1})(2N-1-\alpha_{1})-(2N-1)\geq 2N-1-2\alpha_{1}>0$ and hence $(i_{2}-i_{1})\alpha_{1}<(2N-1)(i_{4}-i_{3})$. Now it follows from Equation 6 that $(i_{2}-i_{1})\alpha_{1}=\Delta^{\prime\prime}$, where $\Delta^{\prime\prime}=\Delta^{\prime}\pmod{(2N-1)(i_{4}-i_{3})}$. From this, we get $\alpha_{1}=\frac{\Delta^{\prime\prime}}{(i_{2}-i_{1})}$. Then by Equation 5, we have $\alpha_{2}=\frac{\Delta^{\prime}-(i_{2}-i_{1})\alpha_{1}}{(2N-1)(i_{4}-i_{3})}$. Now the proof is completed. ∎
@@ III Bounds
We could not find any existing upper or lower bounds on the maximum size of a $2$-$b$-burst-deletion code. In this section, we will derive these bounds.
Let $M_{q,n,b}$ be the maximum size of a $2$-$b$-burst-deletion correcting code in $\Sigma_{q}^{n}$, where $n>2b$.
@@
$M_{n,q,b}\geq\frac{q^{n-2b}}{\binom{n}{2}^{2}}$.
@@
We construct a graph $G$ where the vertex set $V(G)$ is $\Sigma_{q}^{n}$ and two distinct vertices $\mathbf{x}$, $\mathbf{y}$ in $\Sigma_{q}^{n}$ is connected by an edge (denoted by $\mathbf{x}\sim\mathbf{y}$) if and only if $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})\neq\emptyset$. An independent set of $G$ is a subset of $\Sigma_{q}^{n}$ such that any two distinct vertices are not connected by an edge. Let $\alpha(G)$ be the maximum size of an independent set of $G$. A subset $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a $2$-$b$-burst-deletion correcting code if and only if $\mathcal{C}$ is an independent set in $G$. Therefore, we have $M_{q,n,b}=\alpha(G)$. For a vertex $\mathbf{x}$, let $d(\mathbf{x})$ be the number of $\mathbf{y}$ such that $\mathbf{x}\sim\mathbf{y}$. Then it follows from [38, page 100, Theorem 1] that
@@
In [39, Section IV-B], the authors proved an upper bound of $1$-$b$-burst-deletion correcting codes. Next, we adapt their idea to derive an upper bound on $M_{q,n,b}$. Recall that $\ln(\cdot)$ is the natural logarithm function.
@@
For $q\geq 2$, let $f(q)=\min\left\{\frac{1}{q},\frac{q-1}{2q},\frac{(q-1)^{2}}{q^{2}-3q+6}\left(\frac{1}{q}-\frac{(q-1)\ln q}{2q^{3}}\right)\right\}$. When $n\geq 30$ is sufficiently large such that $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$ and $\left(1-\frac{q}{q-1}\sqrt{\frac{12\log n}{n\log q}}\right)^{2}\left(1-b/n\right)^{2}\geq 2/3$, we have
@@
Define $m=n/b-1$. Let $\mathcal{C}\subseteq\Sigma_{q}^{n}$ be a $2$-$b$-busrt-deletion correcting code. Set $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and $t=\left(1-\frac{1}{q}-\epsilon\right)m$. We partition $\mathcal{C}$ into two disjoint subsets: $\mathcal{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}$, where $\mathcal{C}_{1}=\left\{\mathbf{x}\in\mathcal{C}:r(A(\mathbf{x})_{i})\geq t+2\text{ for some }1\leq i\leq b\right\}$ and $\mathcal{C}_{2}=\left\{\mathbf{x}\in\mathcal{C}:r(A(\mathbf{x})_{i})\leq t+1\text{ for all }1\leq i\leq b\right\}$ (recall that $r(\cdot)$ denotes the number of runs). To derive an upper bound of $\left|\mathcal{C}\right|$, it is sufficient to upper bound $\left|\mathcal{C}_{1}\right|$ and $\left|\mathcal{C}_{2}\right|$.
@@
$\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|\geq\max_{1\leq i\leq b}\left|\mathcal{B}_{2}(A(\mathbf{x})_{i})\right|\geq\binom{\max_{1\leq i\leq b}\left\{r\left(A(\mathbf{x})_{i}\right)\right\}-1}{2}\geq\binom{t+1}{2},$ (8)
for all $\mathbf{x}\in\mathcal{C}_{1}$. Since $\mathcal{C}$ is a $2$-$b$-burst-deletion correcting code, we know that $\mathcal{C}_{1}$ is also a $2$-$b$-burst-deletion correcting code. So we have $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})=\emptyset$ for all distinct $\mathbf{x},\mathbf{y}\in\mathcal{C}_{1}$. Then it follows from Equation 8 that $\left|\mathcal{C}_{1}\right|\binom{t+1}{2}\leq\sum_{\mathbf{x}\in\mathcal{C}_{1}}\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|\leq q^{n-2b}$ and hence
$\left|\mathcal{C}_{1}\right|\leq\frac{q^{n-2b}}{\binom{t+1}{2}}\leq\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2n^{2}}{t^{2}}$ $=\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2n^{2}}{\left(1-1/q-\epsilon\right)^{2}\left(n/b-1\right)^{2}}$ (9)
@@
Next, we proceed to upper bound $\left|\mathcal{C}_{2}\right|$. To that end, define a set
$\mathcal{C}^{\prime}\triangleq\left\{\mathbf{x}=\mathbf{x}^{(1)}\cdots\mathbf{x}^{(b)}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}^{(i)}\in\Sigma_{q}^{n/b},\;\forall 1\leq i\leq b\\ r\left(\mathbf{x}^{(i)}\right)\leq t+1,\;\forall 1\leq i\leq b\end{array}\right\}.$

### v5

@@ II-A Related Results
Recently, there have been two notable developments in non-binary two-deletion correcting codes with low redundancy. In [21, Theorem 1], the authors presented a $q$-ary two-deletion correcting code with redundancy at most $5\log n+(16\log q+10)\log\log n+o(\log\log n)$, for any even $q>2$. In [22], a $q$-ary two-deletion correcting code was constructed, with redundancy at most $5\log n+10\log\log n+O_{q}(1)$ (where $O_{q}(1)$ denotes a constant depending only on $q$), for any $q>2$. In Section V, we present a new construction of $q$-ary two-deletion correcting codes with redundancy at most $5\log n+(14\left\lceil\log q\right\rceil+11)\log\log n+O_{q}(1)$, for any $q>2$.
Regarding codes that can correct two $b$-burst-deletions (where $b>1$), there are, to our knowledge, no explicit results available. A related result can be found in the work of Sima et al [34, Section IV-B], where they considered a more generalized type of burst error pattern: $t$ bursts each of length at most $t_{L}$ with the deletions in each burst not necessarily occurring consecutively. For $t=2$ and $t_{L}=b$, their result provides a binary 22-$b$-burst-deletion correcting code with redundancy at most $8\log n+o(\log n)$. Their result was derived using the syndrome compression technique, which was later extended to syndrome compression with pre-coding in [20]. We will apply this extended technique to provide a $q$-ary code with redundancy at most $7\log n+O\left(\log n/\log\log n\right)$, for all $q\geq 2$.
@@
In other words, $\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)$ is the set of all sequences (except $\mathbf{x}$) in $\mathcal{E}$ whose error-ball intersects with that of $\mathbf{x}$.
@@ 1. $(\textup{{P}}1)$
Then there exists a function $\bar{f}:\mathcal{E}\rightarrow\{0,1\}^{2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)}$, computable in polynomial time [^2] In this paper, when saying that a function is computable in polynomial/linear time, we mean that this function is computable in time polynomial/linear in the code-length $n$. such that $\bar{f}\left(\mathbf{x}\right)\neq\bar{f}\left(\mathbf{y}\right)$ for any $\mathbf{x}\in\mathcal{E}$ and $\mathbf{y}\in\mathcal{N}_{\mathcal{E}}\left(\mathbf{x}\right)$.
Let $\mathcal{E}$ be a 11-$b$-burst correcting code and $\bar{f}$ be given in Lemma II.1. Then Lemma II.1 asserts that if $\mathbf{x},\mathbf{y}\in\mathcal{E}$ are distinct codewords and $\bar{f}(\mathbf{x})=\bar{f}(\mathbf{y})$, we have $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})=\emptyset$. Therefore, for any $\mathbf{a}\in\{0,1\}^{2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)}$, the code $\mathcal{E}^{\prime}=\left\{\mathbf{x}\in\mathcal{E}:\bar{f}(\mathbf{x})=a\right\}$ is a 22-$b$-burst-deletion correcting code. Furthermore, by the pigeonhole principle, there exists an $\mathbf{a}$ such that the redundancy of $\mathcal{E}^{\prime}$ is at most $\rho\left(\mathcal{E}^{\prime}\right)=\rho(\mathcal{E})+2\log(N)+O\left(\frac{R(n)}{\log\left(R(n)\right)}\right)$. [^3] We select $\mathcal{E}$ to be a 11-$b$-burst correcting code to obtain a better redundancy. If we take $\mathcal{E}=\Sigma_{q}^{n}$ we get a redundancy of $8\log n+o(\log n)$.
@@
[32, Theorem 9, $t=b,s=0$] For all $q\geq 2$ and $n\geq b$, there is a function $\phi:\Sigma_{q}^{n}\rightarrow\Sigma_{2}^{\log n+O_{q,b}(1)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{q}^{n}$, given $\phi(\mathbf{x})$ and $\mathbf{y}\in\mathcal{B}_{1}^{b}\left(\mathbf{x}\right)$, one can uniquely and efficiently recover $\mathbf{x}$. Here, $O_{q,b}(1)$ is a constant dependent only on $q$ and $b$.
This lemma gives a 11-$b$-burst-deletion correcting code $\mathcal{E}$ with redundancy $\log n+O_{q,b}(1)$. Since $\mathcal{E}$ can correct single $b$-burst-deletion, by a simple counting, we can see that $\left|\mathcal{N}_{\mathcal{E}}(\mathbf{x})\right|<q^{b}n^{3}$. In fact, each codeword in $\mathcal{N}_{\mathcal{E}}(\mathbf{x})$ can be obtained in the following three steps:
@@
    Delete two substrings of length $b$ from $\mathbf{x}$, resulting in $\mathbf{z}^{(1)}$. There are less than $n^{2}$ possibilities for $\mathbf{z}^{(1)}$.
@@
    Insert a sequence of length $b$ into $\mathbf{z}^{(2)}$ to get a sequence $\mathbf{y}\in\Sigma_{q}^{n}$. Since $\mathcal{E}$ is a 11-$b$-burst-deletion correcting code, for each $\mathbf{z}^{(2)}$, there is at most one $\mathbf{y}$ which is in $\mathcal{N}_{\mathcal{E}}(\mathbf{x})$.
@@
[17, Theorem 2] For any integer $n\geq 3$, there exists a function $\xi:\Sigma_{2}^{n}\rightarrow\Sigma_{2}^{7\log n+O\left(1\right)}$, computable in linear time, such that for any $\mathbf{x}\in\Sigma_{2}^{n}$, given $\xi\left(\mathbf{x}\right)$ and any $\mathbf{y}\in\mathcal{B}_{2}\left(\mathbf{x}\right)$ (i.e., $\mathbf{y}$ is obtained from $\mathbf{x}$ by two deletions), one can uniquely and efficiently recover $\mathbf{x}$.
@@
where $x_{k,i}\in\{0,1\}$ such that $x_{i}=\sum_{k=1}^{\left\lceil\log q\right\rceil}x_{k,i}2^{k-1}$ for all $1\leq i\leq n$.
@@
Let $A(\mathbf{x})=A(\mathbf{x},b)$ and $A(\mathbf{y})=A(\mathbf{y},b)$. Since $\mathbf{y}\in\mathcal{B}_{2}^{b}(\mathbf{x})$, we have $A(\mathbf{y})_{i}\in\mathcal{B}_{2}\left(A(\mathbf{x})_{i}\right)$ for all $1\leq i\leq b$. Let $\xi_{1}(\cdot)$ be the function defined in Lemma II.4. Define $\psi(\mathbf{x})\triangleq\left(\xi_{1}\left(A(\mathbf{x})_{1}\right),\ldots,\xi_{1}\left(A(\mathbf{x})_{b}\right)\right)$. Then by Lemma II.4, given $\mathbf{y}\in\mathcal{B}_{2}^{b}\left(\mathbf{x}\right)$ and $\psi(\mathbf{x})$, one can uniquely and efficiently recover $\mathbf{x}$. Since each $\xi_{1}(A(\mathbf{x})_{i})$ is a binary vector of length $7\left\lceil\log q\right\rceil\log(n/b)+O_{q}(1)$, we can see that $\psi(\mathbf{x})$ is a binary vector of length $7b\left\lceil\log q\right\rceil\log(n/b)+O_{q,b}(1)$. ∎
@@
Before proceeding to subsequent sections, we introduce a useful lemma, which will be used in Section IV-B.
Let $N\geq 2$ be an integer. Suppose $\mathbf{x}\in[0,N-1]^{n}$, where $n>4$. Let ?? denote an unknown symbol (not in [0,N1][0,N-1]). If $\mathbf{y}\in\left([0,N-1]\cup\{?\}\right)^{n}$ such that $y_{i}=y_{i+1}=y_{j}=y_{j+1}=?$ and $y_{k}=x_{k}$ for any $k\notin\{i,i+1,j,j+1\}$, we say that $\mathbf{y}$ is obtained from $\mathbf{x}$ by two bursts of erasures (of length two). For our purpose in Section IV-B, assume $j\geq i+2$. For a sequence $\mathbf{x}$ over the alphabet [0,N1][0,N-1], denote $\textup{Syn}\left(\mathbf{x}\right)\triangleq\sum_{i=1}^{n}ix_{i}$. In the following lemma, denote $A(\mathbf{x})=A(\mathbf{x},2)$, i.e., the matrix representation of $\mathbf{x}$ with two rows.
Lemma II.6
For any $0\leq a_{1},a_{2}<2N$ and $0\leq b<nN^{2}$, define $\mathcal{C}$ to be the set of all sequences $\mathbf{x}\in[0,N-1]^{n}$ that satisfies:
@@
[…]
$\quad\sum_{j=1}^{\left\lceil n/2\right\rceil}A(\mathbf{x})_{2,j}\pmod{2N}$,
$\quad W(\mathbf{x})\pmod{nN^{2}}\Bigg{)}$.
@@
$\Delta=$ $\left(b-\sum_{j=1,j\neq i_{1},i_{2}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{y})_{1,j}\right$. (2)
$\quad\quad\left.-(2N-1)\sum_{j=1,j\neq i_{3},i_{4}}^{\left\lceil n/2\right\rceil}j\cdot A(\mathbf{y})_{2,j}\right)\pmod{nN^{2}}$.
@@
Since each term in the right-hand side of (2) is known, we can obtain the value of $\Delta$. Furthermore, by (2), Condition (C2), and the relationship between $A(\mathbf{x})$ and $A(\mathbf{y})$, we have (3).
@@
Since $i_{1}<i_{2}\leq\left\lceil n/2\right\rceil$ and $i_{4}-i_{3}\leq\left\lceil n/2\right\rceil$, we have $1\leq i_{2}-i_{1},i_{4}-i_{3}\leq n/2$. Combining this with the fact $0\leq\alpha_{1},\alpha_{2}\leq N-1$, we conclude that $0\leq(i_{2}-i_{1})\alpha_{1}+(2N-1)(i_{4}-i_{3})\alpha_{2}<nN^{2}$. Now Equation 4 implies
@@
Note that $0\leq\alpha_{1}\leq N-1$. When $i_{4}-i_{3}=i_{2}-i_{1}$ or $i_{2}-i_{1}+1$, it is easy to see that $0\leq(i_{2}-i_{1})\alpha_{1}<(2N-1)(i_{4}-i_{3})$. When $i_{4}-i_{3}=i_{2}-i_{1}-1$, since $i_{2}-i_{1}\geq 2$, we have $(2N-1)(i_{4}-i_{3})-(i_{2}-i_{1})\alpha_{1}=(i_{2}-i_{1})(2N-1-\alpha_{1})-(2N-1)\geq 2N-1-2\alpha_{1}>0$ and hence $(i_{2}-i_{1})\alpha_{1}<(2N-1)(i_{4}-i_{3})$. Now it follows from Equation 6 that $(i_{2}-i_{1})\alpha_{1}=\Delta^{\prime\prime}$, where $\Delta^{\prime\prime}=\Delta^{\prime}\pmod{(2N-1)(i_{4}-i_{3})}$. From this, we get $\alpha_{1}=\frac{\Delta^{\prime\prime}}{(i_{2}-i_{1})}$. Then by Equation 5, we have $\alpha_{2}=\frac{\Delta^{\prime}-(i_{2}-i_{1})\alpha_{1}}{(2N-1)(i_{4}-i_{3})}$. Now the proof is completed. ∎
@@ III Bounds
We could not find any existing upper or lower bounds on the maximum size of a 22-$b$-burst-deletion code. In this section, we will derive these bounds.
Let $M_{q,n,b}$ be the maximum size of a 22-$b$-burst-deletion correcting code in $\Sigma_{q}^{n}$, where $n>2b$.
@@
The maximum size of a 22-$b$-burst-deletion correcting code satisfies $M_{n,q,b}\geq\frac{q^{n-2b}}{\binom{n}{2}^{2}}$.
@@
We construct a graph $G$ where the vertex set $V(G)$ is $\Sigma_{q}^{n}$ and two distinct vertices $\mathbf{x}$, $\mathbf{y}$ in $\Sigma_{q}^{n}$ is connected by an edge (denoted by $\mathbf{x}\sim\mathbf{y}$) if and only if $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})\neq\emptyset$. An independent set of $G$ is a subset of $\Sigma_{q}^{n}$ such that any two distinct vertices are not connected by an edge. Let $\alpha(G)$ be the maximum size of an independent set of $G$. By definition, a subset $\mathcal{C}\subseteq\Sigma_{q}^{n}$ is a 22-$b$-burst-deletion correcting code if and only if $\mathcal{C}$ is an independent set in $G$. Therefore, we have $M_{q,n,b}=\alpha(G)$. For a vertex $\mathbf{x}$, let $d(\mathbf{x})$ be the number of $\mathbf{y}$ such that $\mathbf{x}\sim\mathbf{y}$. Then it follows from [39, page 100, Theorem 1] that
@@
In [40, Section IV-B], the authors proved an upper bound of 11-$b$-burst-deletion correcting codes. Next, we adapt their idea to derive an upper bound on $M_{q,n,b}$. Recall that $\ln(\cdot)$ is the natural logarithm function.
@@
For $q\geq 2$, let $f(q)=\min\left\{\frac{1}{q},\frac{q-1}{2q},\frac{(q-1)^{2}}{q^{2}-3q+6}\left(\frac{1}{q}-\frac{(q-1)\ln q}{2q^{3}}\right)\right\}$. If $n\geq 30$ is sufficiently large such that $\frac{\log n}{n}<\frac{\log q}{12}f(q)^{2}$ and $\left(1-\frac{q}{q-1}\sqrt{\frac{12\log n}{n\log q}}\right)^{2}\left(1-b/n\right)^{2}\geq 2/3$, the maximum size of a 22-$b$-burst-deletion correcting code satisfies
@@
Define m=n/b1m=n/b-1. Let $\mathcal{C}\subseteq\Sigma_{q}^{n}$ be a 22-$b$-burst-deletion correcting code. Set $\epsilon=\sqrt{\frac{12\log n}{n\log q}}$ and $t=\left(1-\frac{1}{q}-\epsilon\right)m$. We partition $\mathcal{C}$ into two disjoint subsets: $\mathcal{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}$, where $\mathcal{C}_{1}=\left\{\mathbf{x}\in\mathcal{C}:r(A(\mathbf{x})_{i})\geq t+2\text{ for some }1\leq i\leq b\right\}$ and $\mathcal{C}_{2}=\left\{\mathbf{x}\in\mathcal{C}:r(A(\mathbf{x})_{i})\leq t+1\text{ for all }1\leq i\leq b\right\}$ (recall that $r(\cdot)$ denotes the number of runs). To derive an upper bound of $\left|\mathcal{C}\right|$, it is sufficient to upper bound $\left|\mathcal{C}_{1}\right|$ and $\left|\mathcal{C}_{2}\right|$.
@@
$\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|$ $\geq\max_{1\leq i\leq b}\left|\mathcal{B}_{2}(A(\mathbf{x})_{i})\right|$ (8)
$\geq\binom{\max_{1\leq i\leq b}\left\{r\left(A(\mathbf{x})_{i}\right)\right\}-1}{2}\geq\binom{t+1}{2}$,
for all $\mathbf{x}\in\mathcal{C}_{1}$. Since $\mathcal{C}$ is a 22-$b$-burst-deletion correcting code, we know that $\mathcal{C}_{1}$ is also a 22-$b$-burst-deletion correcting code. So we have $\mathcal{B}_{2}^{b}(\mathbf{x})\cap\mathcal{B}_{2}^{b}(\mathbf{y})=\emptyset$ for all distinct $\mathbf{x},\mathbf{y}\in\mathcal{C}_{1}$. Then it follows from Equation 8 that $\left|\mathcal{C}_{1}\right|\binom{t+1}{2}\leq\sum_{\mathbf{x}\in\mathcal{C}_{1}}\left|\mathcal{B}_{2}^{b}(\mathbf{x})\right|\leq q^{n-2b}$ and hence
$\left|\mathcal{C}_{1}\right|\leq\frac{q^{n-2b}}{\binom{t+1}{2}}$ $\leq\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2n^{2}}{t^{2}}$ (9)
$=\frac{q^{n}}{n^{2}q^{2b}}\cdot\frac{2n^{2}}{\left(1-1/q-\epsilon\right)^{2}\left(n/b-1\right)^{2}}$
@@
Next, we proceed to upper bound $\left|\mathcal{C}_{2}\right|$. To that end, define $\mathcal{C}^{\prime}$ to be the following set
$\left\{\mathbf{x}=\mathbf{x}^{(1)}\cdots\mathbf{x}^{(b)}\in\Sigma_{q}^{n}:\begin{array}[]{c}\mathbf{x}^{(i)}\in\Sigma_{q}^{n/b},\;\forall 1\leq i\leq b\\ r\left(\mathbf{x}^{(i)}\right)\leq t+1,\;\forall 1\leq i\leq b\end{array}\right\}.$